OpenStudy (anonymous):

if f(x) be a polynomial function satisfying f(x).f(1/x) =f(x) + f(1/x) and f(4)=65 then find f(6)

OpenStudy (asnaseer):

the polynomial can be written down as:\[f(x)=a_0+a_1x+a_2x^2+a_3x^2+...\]and\[f(\frac{1}{x})=a_0+\frac{a_1}{x}+\frac{a_2}{x^2}+\frac{a_3}{x^3}+...\]we are given:\[f(x).f(\frac{1}{x})=f(x)+f(\frac{1}{x})\]therefore:\[f(\frac{1}{x})=\frac{f(x)}{f(x)-1}\]we are given f(4)=65, so:\[f(\frac{1}{4})=\frac{65}{65-1}=\frac{65}{64}\]and\[f(4)=65=a_0+4a_1+16a_2+64a_3+...\]\[f(\frac{1}{4})=\frac{65}{64}=a_0+\frac{a_1}{4}+\frac{a_2}{16}+\frac{a_3}{64}+...\]therefore, multiplying both sides of the last equation by 64, we get:\[65=64a_0+16a_1+4a_2+a_3+...\] that's as far as I have been able to take this. is there any other information given with this question or are you able to continue from here?

OpenStudy (anonymous):

i am trying from here....thanks

OpenStudy (asnaseer):

yw - do let me know if you happen to solve it - I'd be very interested in the solution.

OpenStudy (mr.math):

I think I found a function that satisfies the given conditions. Take \(f(x)=x^2+1\)!

OpenStudy (asnaseer):

is that through trial and error or analysis?

OpenStudy (asnaseer):

but f(4) != 65?

OpenStudy (mr.math):

I meant f(x)=x^3+1 :)

OpenStudy (asnaseer):

f(x)=1+16x will also work, as will many others (I think) but which one is correct?

OpenStudy (mr.math):

Actually any polynomial of the form \(x^n+1\), satisfies \(f(x)f(\frac{1}{x})=f(x)+f(\frac{1}{x})\).

OpenStudy (mr.math):

Right?

OpenStudy (asnaseer):

I don't think so - e.g. \(1+x^3\) does not work

OpenStudy (asnaseer):

I mean \(1+x\)

OpenStudy (asnaseer):

there must be another condition to pin down the solution to just one polynomial

OpenStudy (mr.math):

(1+x)(1/x+1)=1/x+1+1+x, right?

OpenStudy (asnaseer):

yes but it must also fit f(4)=65

OpenStudy (mr.math):

I was talking about the statement I said about polynomial of the form x^n+1. x^3+1 satisfies the conditions in the problem, but as you said there are many other polynomials that do, and they would give different values for f(6). So, I think there's something missing in the problem.

OpenStudy (asnaseer):

actually Mr.Math I think you may have something there with \(x^n+1\)

OpenStudy (anonymous):

note that if you plug in x = 1, you get:\[f(1)^2 = 2f(1)\]which says f(1) = 0 or f(1) = 2. not that i know where to go, but that may help.

OpenStudy (mr.math):

Wait! Could you give me another example of such a polynomial?

OpenStudy (mr.math):

If this was in an exam, I would say \(f(6)=6^3+1\). :-)

OpenStudy (asnaseer):

I think you are right

OpenStudy (asnaseer):

because with f(x)=16x+1, we get f(4)=65, but it does not satisfy the original conditions.

OpenStudy (mr.math):

Thank you! :-)

OpenStudy (asnaseer):

Although I would be a lot happier if we could prove this

OpenStudy (mr.math):

So am I.

OpenStudy (asnaseer):

i.e. prove that there are no other polynomials that could satisfy the original condition.

OpenStudy (mr.math):

I don't think that we have to do that. The question doesn't say that in anyway. We have only to show that f(6) of any such polynomial is 6^3+1.

OpenStudy (mr.math):

any way*

OpenStudy (mr.math):

Do you agree?

OpenStudy (asnaseer):

I think you have found the answer, but I personally would be interested in a proof that this is indeed the only such polynomial (which I believe it must be for the question to make any sense).

OpenStudy (asnaseer):

ok, I think I have a proof. from above I proved that:\[f(\frac{1}{x})=\frac{f(x)}{f(x)-1}\]and we can write:\[\begin{align} f(\frac{1}{x})=a_0+\frac{a_1}{x}+\frac{a_2}{x^2}+\frac{a_3}{x^3}+...+\frac{a_n}{x^n}\\ =\frac{1}{x^n}(a_0x^n+a_1x^{n-1}+a_2x^{n-2}+a_3x^{n-3}+...+a_n)\\ =\frac{f(x)}{f(x)-1}\quad\text{from above}\\ \therefore (f(x)-1)(a_0x^n+a_1x^{n-1}+a_2x^{n-2}+a_3x^{n-3}+...+a_n)=x^nf(x)\\ \therefore (a_0-1+a_1x+...+a_nx^n)(a_0x^n+a_1x^{n-1}+...+a_n)=x^n(a_0+a_1x+...+a_nx^n)\\ =a_0x^n+a_1x^{n+1}+...+a_nx^{2n} \end{align}\] from inspection, we can infer:\[a_n(a_0-1)=0\implies a_n=0\quad\text{or }a_0=1\] 1. if we pursue the case of \(a_n=0\) then we will successively get \(a_{n-1}=0\), \(a_{n-2}=0\), ..., \(a_1=0\) and will end up with \(a_0=2\). this gives us \(f(x)=f(\frac{1}{x})=2\) and satisfies the condition \(f(x).f(\frac{1}{x})=f(x)+f(\frac{1}{x})\) but does not satisfy the condition \(f(4)=65\). so we can reject this case. 2. if we pursue the case \(a_0=1\) we will successively be able to prove that \(a_1=0\), \(a_2=0\), ...\(a_{n-1}=0\), and will end up with:\[a_nx^n(x^n+a_n)-x^n+a_nx^{2n}\]from which we can prove \(a_n=1\). this gives us \(f(x)=1+x^n\) and satisfies the condition \(f(x).f(\frac{1}{x})=f(x)+f(\frac{1}{x})\). in order to satisfy the secondary condition \(f(4)=65\), we need \(n=3\). so we have proved that the only solution is:\[f(x)=1+x^3\]therefore:\[f(6)=1+6^3=1+216=217\] In conclusion - you were right all along Mr.Math :-)

OpenStudy (mr.math):

Great! Nice work. If I might add that even the case where f(x)=2 is of the form 1+x^n, with n=0. Thus we can conclude (based on your proof) that ONLY polynomials of the form 1+x^n (n>=0) satisfies that condition.

OpenStudy (asnaseer):

good point - well spotted and thx for your insights :-)

OpenStudy (anonymous):

thanks guys..